what is the length of a base of a traingle with a height of 864 and a area of 1871
Answer by checkley77(12844)

AREA=BH/2
1,871=B*864/2
1,871=864B/2
864B=1,871*2
864B=3,742
B=3,742/864
B=4.331 ANS. FOR THE BASE MEASUREMENT.
PROOF:
1,871=4.331*864/2
1,871=3,742//2
1,871=1,871